# Nuclear cross sections from low-energy interactions

### v1

@@ III Results
III.1 ${}^{24}\mathrm{Mg}+n$ scattering
As the first implementation of this method, we have calculated total and elastic neutron scattering cross sections of the characteristically prolate deformed nucleus ${}^{24}\mathrm{Mg}$.
Following the procedure described in [11], the effective Hamiltonian was created using the SLy4 Skyrme parameterization. The number of major shells in the harmonic oscillator single–particle basis was 17 in the HF calculation, but the $586$ lowest energy HF states (which correspond to approximately as many states as there would be in 11 major shells) were used as the single–particle basis for the GCM. We then sampled the $\beta\gamma$ plane with 275 states. Keeping states below $25\leavevmode\nobreak\ \mathrm{MeV}$ resulted in 179 HFB basis states. The states were projected with 10 particle number projection points, and $(6,12,24)$ angular momentum projection points for the $(\alpha,\beta,\gamma)$ angles and $b=0.45$. The odd–even HFB basis consisted of one quasiparticle excitation for each even–even HFB state. The quasiparticle to excite was chosen randomly among the 10 quasiparticles with the lowest mean-field energy of a given signature and $J\pi$ such that quasiparticle states that correspond to the wrong particle number are also excluded, i.e. hole states when calculating the solution for $A+1$. These parameters were chosen to ensure optimal convergence in the energy range presented. Using this Hamiltonian and basis states, the states of ${}^{23,24,25}\mathrm{Mg}$ were calculated [11, 18]. The wavefunctions were then used to calculate the spectroscopic amplitudes of ${}^{23}\mathrm{Mg}$ and ${}^{25}\mathrm{Mg}$ relative to the ground state of ${}^{24}\mathrm{Mg}$.
Then, for the scattering calculations, $\eta$ was set to $\eta=\frac{a}{\pi}\frac{\left(E-E_{f}\right)^{2}}{\left(E-E_{f}\right)^{2}+b^{2}}$ with $a=12\leavevmode\nobreak\ \mathrm{MeV}$ and $b=22.36\leavevmode\nobreak\ \mathrm{MeV}$ as in [26].
@@ III.1 ${}^{24}\mathrm{Mg}+n$ scattering
Fig. 1: The total scattering cross section of ${}^{24}\mathrm{Mg}+n$ in function of the neutron energy. The red and purple lines correspond to the cross sections calculated using the method described in this paper with signature $+i$ and $-i$ respectivey and the area between them is shaded red. The dashed line are result using the Koning–Delaroche optical potential. The black circles are experimental cross sections for ${}^{24}\mathrm{Mg}$ [27], while the diamonds are natural Magnesium [28].
The cross sections are calculated for the two signatures ±$\pm i$. The signature is a preserved symmetry and should not have effect on observables in an exact calculation, but due to the random selection of basis states and quasiparticle excitation in the GCM, and the inclusion of cranking, quasiparticle states of different signature behave differently [29]. Comparing the two can give an indication of the convergence of the GCM calculation.
@@
In Fig. 1 the integrated neutron total cross sections are shown for energies from $E=0.5\leavevmode\nobreak\ \mathrm{MeV}$ to $13\leavevmode\nobreak\ \mathrm{MeV}$, together with the result of Koning–Delaroche optical potential [5]. We can see a good agreement with experimental data for the whole region, and the phenomenological Koning–Delaroche results lie in the middle of our predicted interval for most of the region below 10 MeV. Due to the choice of $\eta$, resonances are wider than what experiments show at low energies.
@@
Fig. 2: The elastic scattering cross section of ${}^{24}\mathrm{Mg}+n$ in function of neutron energy. The lines and datapoints mean the same as in Fig. 1. Additionally, the integrated cross section calculated with only the HF potential is shown as a dashed blue line. The experimental cross sections for ${}^{24}\mathrm{Mg}$ are from [30, 31, 32], while the natural Magnesium cross sections are from [33, 34, 35, 36, 37, 38, 39, 40].
@@
The elastic scattering cross section is also calculated and compared to experiment in Fig. 2. We see that the calculated cross section reproduces experiments up to $9\leavevmode\nobreak\ \mathrm{MeV}$, but overestimates them for neutron energies above $10\leavevmode\nobreak\ \mathrm{MeV}$. This indicates that the method fails to find enough states that contribute to the non–elastic channels above $9\leavevmode\nobreak\ \mathrm{MeV}$, and progressively more of the contribution comes from the HF solution through the completion. The static HF potential is real and only the elastic scattering channel is available in this case, resulting in a sizeable overestimation of the elastic scattering cross section in Fig. 2. At higher energy the calculated GCM $\sigma_{E,T}$ will gradually reach the HF result above 20 MeV.
@@
Fig. 3: The differential cross section of elastic ${}^{24}\mathrm{Mg}+n$ scattering for four different neutron energies. The energies are shown in the figure in MeV. The lines and datapoints mean the same as in Fig. 2. The dashed line refer to the HF calculation at 1.5 and 9.76 MeV. The experimental data are from [35, 31, 41, 30]. The $1.5\leavevmode\nobreak\ \mathrm{MeV}$ reaction is represented in laboratory frame, while the other energies are represented in center of mass frame. The cross sections for each energy are shifted down by a factor of 10 relative to the previous.
The angular differential cross sections were calculated for several projectile energies shown in Fig. 3, together with the corresponding experimental values. We see also for the differential cross section a general agreement with experiment. For some energies this method underestimates the cross section at the minima, which could be due to an underestimation of the compound nucleus reaction channels. This is when the neutron is absorbed and an excited ${}^{25}\mathrm{Mg}$ is formed. This compound nucleus will then decay and one possible decay mode is to emit a neutron. If the emitted neutron has the same energy as the original incoming neutron, the result is identical to elastic scattering. Since the compound nucleus has time to thermalize, the emission is isotropic and it will contribute an angle independent term to the differential cross section. The compound nucleus contribution is expected to be larger at lower energies, and in Fig. 3 the results at lower neutron energies can be seen to underestimate the minima in the experimental cross sections. The effect of the compound nucleus are particularly challenging to model from a Hamiltonian and is under further investigation [42].
@@
In conclusion, we have found that the lack of absorption observed when using a microscopically generated optical potential, that was linked to configurations beyond particle–hole excitation in [14], is properly taken into account through the collective degrees of freedom when using GCM. This shows that GCM captures the necessary low energy correlations needed to accurately describe low energy neutron scattering to a level that was not previously achieved. This work extends microscopic reaction approaches to study deformed, heavy, and exotic nuclei, representing a significant step towards a unified model of nuclear structure and reaction.

### v2

@@ III Results
As the first implementation of this method, we have calculated total and elastic neutron scattering cross sections of the characteristically prolate deformed nucleus $^{24}\mathrm{Mg}$. The calculation of the nuclear wavefunctions were executed following the framework described in [12, 27]. The effective Hamiltonian was created for 24Mg using the SLy4 Skyrme parameterization [32]. SLy4 was previously used to calculate properties of 24Mg in [12, 33]. Using this Hamiltonian and basis states, the eigenstates of $^{23,24,25}\mathrm{Mg}$ were calculated [12, 27]. The convergence parameters are described in [21]. The wavefunctions were then used to calculate the spectroscopic amplitudes of $^{23}\mathrm{Mg}$ and $^{25}\mathrm{Mg}$ relative to the ground state of $^{24}\mathrm{Mg}$ and to construct the corresponding Green’s functions (7).
For the scattering calculations, $\eta$ was set to $\eta=\frac{a}{\pi}\frac{\left(E-E_{f}\right)^{2}}{\left(E-E_{f}\right)^{2}+b^{2}}$ with $a=12\leavevmode\nobreak\ \mathrm{MeV}$ and $b=22.36\leavevmode\nobreak\ \mathrm{MeV}$ following a common prescription to represent the average resonance widths [34, 35]. A variation of a factor 4 of the $\eta$ parameter obtains scattering results generally close to the provided convergence band, as shown in the inset of Fig. 1. $a=24\leavevmode\nobreak\ \mathrm{MeV}$ gives on average $8.8\%$ lower total cross section than $a=6\leavevmode\nobreak\ \mathrm{MeV}$, much smaller than the factor 4 increase of the average resonance widths represented by $\eta$. Methods for treating the imaginary part of the Green’s function in (7) have been developed for example in [36] using self–consistency or in [1, 23] using continuum states. Such a consistent approach is not within the scope of this paper, but since the method can be seen to perform well for a wide range of $\eta$, it would not change the conclusions of this paper. Neutron scattering on 24Mg was also calculated in [36, 37] with single–reference configuration interaction (shell model) prescriptions, making it an interesting case of comparison.
@@
Fig. 1: The total scattering cross section of $^{24}\mathrm{Mg}+n$ in function of the neutron energy in the lab frame. In the main figure, the red and purple lines correspond to the cross sections calculated using the method described in this paper with signature +i+i and i-i respectively and the area between them is shaded red. The cyan dashdotted line shows the result using the Koning–Delaroche optical potential. The black circles are experimental cross sections for $^{24}\mathrm{Mg}$ [38], while the diamonds are natural Magnesium [39]. In the inset, the results for +i+i signature and variation of $\eta$ parameters, with $a=6,12,24$ MeV for light blue, green and purple line respectively.
The cross sections are calculated using two different odd nucleus bases. These basis states are generated as eigenstates of the intrinsic $\exp(-i\pi j_{x})$ operator with eigenvalue either +i+i or i-i known as their signature quantum number. Since either basis forms a complete set of many-body states after angular momentum projection [15], comparing the two can give an indication of the convergence of the cross section calculation. Additionally, cross sections were calculated using only the completion states by setting $s^{x}_{i,\alpha}=0$ in (5,6) and are presented in Fig. 2 and 3.
@@
In Fig. 1 the integrated neutron total cross sections are shown for energies from $E=0.5\leavevmode\nobreak\ \mathrm{MeV}$ to $13\leavevmode\nobreak\ \mathrm{MeV}$, together with the result of Koning–Delaroche optical potential [6]. We can see that below 10 MeV the calculation overestimates the cross section. Due to the choice of $\eta$, resonances are wider than what experiments show at energies below $5\leavevmode\nobreak\ \mathrm{MeV}$.
@@
Fig. 2: The elastic scattering cross section of $^{24}\mathrm{Mg}+n$ in function of neutron energy in the lab frame. The lines and datapoints mean the same as in Fig. 1. Additionally, the integrated cross section calculated with only the completion states is shown as a dashed blue line. The experimental cross sections for $^{24}\mathrm{Mg}$ are from [40, 41, 42], while the natural Magnesium cross sections are from [43, 44, 45, 46, 47, 48, 49, 50].
@@
The elastic scattering cross section is also calculated and compared to experiment in Fig. 2. We see that the calculated cross section slightly overestimates the cross section up to $9\leavevmode\nobreak\ \mathrm{MeV}$, and this difference increases for neutron energies above $10\leavevmode\nobreak\ \mathrm{MeV}$. This indicates that the method fails to find enough states that contribute to the non–elastic channels above $9\leavevmode\nobreak\ \mathrm{MeV}$, and progressively more of the contribution comes from the mean field through the completion. The potential obtained using only completion states is a real potential and only the elastic scattering channel is available in this case, resulting in a sizeable overestimation of the elastic scattering cross section in Fig. 2. At higher energy the calculated GCM $\sigma_{E,T}$ will gradually reach the result with only completion above 20 MeV.
@@
Fig. 3: The differential cross section of elastic $^{24}\mathrm{Mg}+n$ scattering for four different neutron energies with respect to center–of–mass angle. The lab frame energies are shown in the figure in MeV. The lines and datapoints mean the same as in Fig. 2. The dashed lines at 1.5 and 9.76 MeV refer to the calculation with only the completion states. The experimental data are from [45, 41, 51, 40]. The cross sections for each energy are shifted down by a factor of 10 relative to the previous.
The angular differential cross sections were calculated for several projectile energies, as shown in Fig. 3, along with the corresponding experimental values. It is important to note that the solution to the scattering problem represents only one component of the experimentally measured cross section. When a neutron is absorbed, forming an excited $^{25}\mathrm{Mg}$ nucleus that subsequently decays by emitting another neutron of the same energy, this compound nucleus reaction contributes to the measured cross section and is indistinguishable from elastic scattering. Since the compound nucleus has time to thermalize, its emission is approximately isotropic, contributing to the differential cross section with little angular dependence. This effect is expected to be more significant at lower energies and around the minima.
To more accurately compare with experimental data, in Fig. 3 we include the compound nucleus contribution calculated using the semi-microscopic optical model of [52, 53] (known as JLM), which is based on properties of nuclear matter and it is phenomenologically adjusted to reproduce scattering. Developing a compound nucleus model for a general non–local potential is particularly challenging and beyond the scope of the present study, hence we use a local optical model as in [36]. This additional component is necessary because of the use of scattering theory (8) to describe elastic scattering, not due to our particular microscopic approach constructing the optical potential.
@@
In conclusion, an energy weighted sum rule for spectroscopic amplitudes is introduced, interpreted as a generalized mean field, which bridges the gap between nuclear structure and scattering that arise due to the necessary truncations done in many–body methods. The results demonstrate the viability of this new method, and it enables calculating scattering observables from a wide range of many–body methods and interactions. In addition, the results suggest the lack of absorption observed when using a microscopically generated optical potential, that was linked to configurations beyond particle–hole excitation in [16], is partially taken into account through the collective degrees of freedom when using GCM. This shows that GCM can capture necessary low energy correlations needed to describe low energy neutron scattering with a microscopic method using multiple reference states, though further research is required to better understand the resonance widths of GCM states, important in the description of both elastic and nonelastic scattering. A need for compound scattering implementations for general non–local potentials was also identified. This work extends microscopic reaction approaches to study deformed, heavy, and exotic nuclei, representing a significant step towards a unified model of nuclear structure and reaction. The flexibility of this method will allow for further improvements through the use of new interactions, functionals, collective coordinates, and many–body expansions, and enables it to be used for the whole nuclear chart.